OpenStudy (anonymous):

Pleaseeeeee help!!!:) medal reward! Pleaseeeee help One-sixth of a number is 5 more than the same number. What's the number?

OpenStudy (anonymous):

1/6=5+5 Idk

OpenStudy (anonymous):

1/6 x=5+x

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

The x is not part of the denominator of 1/6, just adding, and your welcome :D

OpenStudy (anonymous):

K

OpenStudy (anonymous):

I know the answers 6 but I just don't get how to get that

OpenStudy (anonymous):

Not 6. "One-sixth of a number is 5 more than the same number" is written mathematically as: 1/6 x = x + 5 Start by multiplying all terms by 6: 6/6 x = 6x + 30, which is the same as: x = 6x + 30 Isolate the x terms: x - 6x = 30 yields: -5x = 30 Divide both sides by -5: x = -6